OpenStudy (darkbeautystar):

Help!! I need someone to explain to me how to do this!!! Misha wants to know how many pepperonis to put on his pizza. He was told that he should use 1 pepperoni rounded to the nearest whole pepperoni for each .75 inches of diameter in a pizza. Show an equation that would solve how many pepperonis he would need if his pizza has a 8 inch radius.

OpenStudy (darkbeautystar):

Are you good at math!?

OpenStudy (anonymous):

I'm pretty good

OpenStudy (darkbeautystar):

Ok...well I need you or someone to draw me the equation it is asking for while explaining what you are doing!!!

OpenStudy (darkbeautystar):

I have to show my work...and I don't know how!

OpenStudy (darkbeautystar):

I need the equation done first.... and then if you explain to me how to do it...I will show my work!

OpenStudy (anonymous):

Ok, I know how to do this, but I'm not sure how to explain it, since it's more of a logical application of a concept.

OpenStudy (darkbeautystar):

Ok.. I will have some else come look at it!!

OpenStudy (darkbeautystar):

@Dan815 -- If you are good in math.. I need your help!!!

OpenStudy (darkbeautystar):

@Hero - Will you please come help me with one math question!?

OpenStudy (darkbeautystar):

@sleepyjess -- Will you please come help me with math? Only one question!

OpenStudy (darkbeautystar):

@SolomonZelman -- Would you come help me with one math question!?

OpenStudy (darkbeautystar):

@paki -- If you are good in math.. would you come help me!? I have one question to do!

OpenStudy (darkbeautystar):

BRB

OpenStudy (darkbeautystar):

@mathstudent55 -- would you come help me with 1 math question!!

OpenStudy (darkbeautystar):

@zepdrix -- would you come help me with 1 math question!?

OpenStudy (darkbeautystar):

@AlexandervonHumboldt2 -- would you come help me with 1 math question!?

OpenStudy (darkbeautystar):

Hi!

zepdrix (zepdrix):

ey c:

zepdrix (zepdrix):

What are we doing? making pizza or something?

OpenStudy (darkbeautystar):

Well... I have this one math question to do.. it has to be done before the 13th...so I am doing it today!

OpenStudy (xmissalycatx):

1 Pepperoni per every .75 inches out of 8 inches..

OpenStudy (darkbeautystar):

So I need the equation written out that they are asking for.. and... I need the work shown!

zepdrix (zepdrix):

`1 pepperoni` per `0.75 diameter` should be equivalent to `how many pepperonis` per `16 diameter` Maybe we'll call the unknown number of pepperonis P,\[\large\rm \frac{1pep}{0.75dia}=\frac{P~pep}{16dia}\] Do you understand why I used 16 instead of 8? :)

OpenStudy (xmissalycatx):

So you would have a TOTAL of an 8 inch radius (which means 16 inches diameter) so that is \[\frac{ 1 }{ .75 } = \frac{ x }{ 16 }\]

OpenStudy (xmissalycatx):

So you would cross multiply which gives you .75x = 16 then you'd solve for x..

zepdrix (zepdrix):

I don't understand why you copy and pasted everything I just did.... ugh

OpenStudy (xmissalycatx):

I didn't copy and paste. I was typing before you had been but you sent it first. My computer was running slow..

OpenStudy (darkbeautystar):

I kinda of understand why you used 16!!

OpenStudy (darkbeautystar):

So my equation would be...1/.75 = x/16

OpenStudy (xmissalycatx):

You use 16 because it gave you your RADIUS (which was 8). A radius is always HALF of the diameter, so you multiply 8 by 2.

OpenStudy (darkbeautystar):

Is that how I would write it out to my teacher!?

OpenStudy (xmissalycatx):

Your equation is .75x = 16

OpenStudy (darkbeautystar):

And then how would I show my work!?

OpenStudy (xmissalycatx):

The 1/.75 = x/16 is how you get your equation..

OpenStudy (xmissalycatx):

You show your work by dividing both sides by .75..

OpenStudy (darkbeautystar):

So I would say .75x=16 .75x/.75x=16/.75

OpenStudy (xmissalycatx):